OpenStudy (anonymous):

what maximum height will be reached by a stone thrown straight up with an initial speed of 35 m/s?

OpenStudy (anonymous):

would initial speed equal v0x?

OpenStudy (anonymous):

any tips or hints?

OpenStudy (anonymous):

feels like this is a y equation so it is getting thrown up

OpenStudy (anonymous):

i think since we're dealing with the y-components \(v_iy= 35 m/s\ and\ use\ V_f=0 \ m/s\) because we're looking for the maximum height Viy=initial velocity Vfy= final vertical velocity

OpenStudy (anonymous):

so voy=35?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

0=35-9.8t

OpenStudy (anonymous):

t=3.57

OpenStudy (anonymous):

wait,time is not given, where did you get t=3.57?

OpenStudy (anonymous):

well

OpenStudy (anonymous):

from the equation Vy=Voy-gt

OpenStudy (anonymous):

essentially 0=35-9.8t right?

OpenStudy (anonymous):

you can use the formula \(V_f^2=V_i^2+2ad\), it's more convenient (^_^) yes, that's right also

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

maximum height=y right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait a minute, the answer is 62

OpenStudy (anonymous):

62.5m

OpenStudy (anonymous):

but wait

OpenStudy (anonymous):

I have y=yo+voyt-1/2gt squared

OpenStudy (anonymous):

which is y=o+35(3.57)-1/2(9.8)(3.57)squared

OpenStudy (anonymous):

or am I messing up somewhere?

OpenStudy (anonymous):

t=3.57, right?

OpenStudy (anonymous):

voy=35?

OpenStudy (anonymous):

it looks right to me yes, what's wrong? O.o

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

the answer is 62.5

OpenStudy (anonymous):

but y=35(3.57)-1/2(9.8)(3.57)squared equals 62.5...

OpenStudy (anonymous):

lol I made a mistake

OpenStudy (anonymous):

DERP

OpenStudy (anonymous):

yeah, i tried doing you're method and i got 62.49999 which is approx 62.5 also

OpenStudy (anonymous):

*your

OpenStudy (anonymous):

next problem?

OpenStudy (anonymous):

how many more? O.O

OpenStudy (anonymous):

till 10 right?

OpenStudy (anonymous):

it is 9.24 P.M. at my place

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay next one then XD